OpenStudy (anonymous):

A ball is thrown upward with an initial velocity of 48 ft/sec from a height 640ft. It's height h, in feet, after t seconds is given by h(t)= -16t^2 +48t+640. After how long will the ball reach the ground? The ball will reach the ground in ___seconds

OpenStudy (anonymous):

Can reduce the equation to a simpler equivalent system by dividing all the terms by -16 first, then use quadratic formula (if it's not easily factored). You'll get two solutions, so throw out the one that doesn't make sense (often the negative value.)

OpenStudy (anonymous):

From inspection, it looks factorable, so quadratic formula would be overkill.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Did you find the solution?

OpenStudy (anonymous):

No sir

OpenStudy (anonymous):

8 sec

OpenStudy (anonymous):

\[-16t^2 +48t+640 = (-16)(t^2 -3t-40)\] You're solving for h=0, so you can throw out the (-16) factor. \[t^2 -3t-40 = (t+5)(t-8).\] You can throw out the (t+5) factor because we're not interested in what happened in the past. t-8=0 at t=8s.